OpenStudy (anonymous):

HELP WILL FAN/METAL 5t+8=43 explain you answer

OpenStudy (thewalkingdeadrick):

whats the question im pretty good in math

OpenStudy (anonymous):

inverse operations

OpenStudy (anonymous):

5t+8=43 subtract 8 from both sides 5t=35 divide both sides by 5 t=7

JoelTheBoss (joel_the_boss):

1 Subtract 8 from both sides 5t=43−8 2 Simplify 43−8 to 35 5t=35 3 Divide both sides by 5 t=355 4 Simplify 355 to 7 t=7

OpenStudy (thewalkingdeadrick):



OpenStudy (thewalkingdeadrick):

lol

JoelTheBoss (joel_the_boss):

step 4 says \[\frac{ 35 }{ 5 }\]

OpenStudy (anonymous):

I have another on